OpenStudy (anonymous):

4+7+509-23÷64×3+63-3+825= ?

OpenStudy (callisto):

4+7+509-23÷64×3+63-3+825 =4+7+509 - (23÷64×3 ) +63-3+825 = 4+7+509 +63-3+825 - (23÷64×3 ) = Can you use your calculator to do it?

OpenStudy (anonymous):

Use pemdas to help you. [Parenthesis Exponent Multiplication Division Addition Subtraction]